ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

7. Lie Groups and Lie Algebras

1991 MS Classification: 22E55 Heiermann, Volker, Humboldt University Berlin, Germany A Plancherel formula for the elements of the Hecke algebra of a reductive p-adic group We show a Paley–Wiener theorem for the elements of the Hecke-algebra of the reductive p-adic group. The proof is based on an analogue of the Plancherel-formula of Harish–Chandra.